### v2

# Code Readability in the Age of Large Language Models: An Industrial Case Study from Atlassian

Abstract.
Software engineers spend a significant amount of time reading code during the software development process. This trend is amplified by the emergence of large language models (LLMs) that automatically generate code. However, little is known about the readability of the LLM-generated code and whether it is still important from practitioners’ perspectives in this new era. In this paper, we conduct a survey to explore the practitioners’ perspectives on code readability in the age of LLMs and investigate the readability of our LLM-based software development agents framework, HULA, by comparing its generated code with human-written code in real-world scenarios. Overall, the findings underscore that (1) readability remains a critical aspect of software development; (2) the readability of our LLM-generated code is comparable to human-written code, fostering the establishment of appropriate trust and driving the broad adoption of our LLM-powered software development platform.
Code Readability, Large Language Models, Practitioner’s Perceptions
booktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norway
1. Introduction
Code readability has long been a key focus in software engineering research (tashtoush2013impact; coleman2018aesthetics; fakhoury2018effect; mannan2018towards; alawad2019empirical; johnson2019empirical). For example, Minelli et al. found that programmers dedicate approximately 70% of their time to reading and understanding source code (minelli2015know). Similarly, as Guido Van Rossum, the creator of Python, has noted: “Code is read more often than it is written”(pepeight2001). Thus, writing readable code not only facilitates better comprehension of its functionality (johnson2019empirical) but also significantly reduces software maintenance costs.
In the age of large language models (LLMs) for code, these models have demonstrated remarkable results across various software development tasks (zhao2023survey) (e.g., code completion (takerngsaksiri2024syntax; takerngsaksiri2024student), test case generation (takerngsaksiri2024tdd; alagarsamy2024a3test), code review automation (li2022codereviewer; pornprasit2024gpt; thongtanunam2022autotransform; hong2022commentfinder), vulnerability detection and repair (fu2022linevul; fu2024aibughunter; fu2022vulrepair)). Recently, there has been a growing interest in integrating LLMs into modern Integrated Development Environments (IDEs) (e.g., GitHub Copilot (copilot2024)) to assist software engineers in implementing new features, updating patches, improving code quality, fixing bugs, and resolving software development tasks.
At Atlassian, code readability has always been a cornerstone of software development, as it plays a critical role in facilitating software evolution and maintenance. In the age of large language models (LLMs), like those integrated into Jira (a task management platform) (jiraai2024), the importance of code readability is amplified. Readable code ensures that teams can efficiently collaborate, debug, and enhance their software, reducing technical debt and long-term costs. Recently, we introduced HULA, a human-in-the-loop software development agents framework powered by large language models (takerngsaksiri2024human). While HULA aims to streamline development tasks, a crucial question arises: does readability retain its value in this LLM-assisted paradigm, and if so, why? Additionally, how readable is the code it generates compared to human-written code? Understanding and ensuring code readability in LLM-generated code is vital for Atlassian, since it directly impacts the productivity and satisfaction of software development teams, enabling teams to work efficiently.
In this paper, we aim to investigate the practitioners’ perceptions on the importance, the challenges, and the state of practice of code readability, and investigate the readability of LLM-generated code and human-written code in the context of enterprise software development tasks. Thus, we address the following two RQs.
Figure 1. The user interface of our HULA (Human-in-the-Loop Software Development Agent) framework, seamlessly integrated into Atlassian JIRA (takerngsaksiri2024human).
  1. RQ1)
    How and why code readability is important in the age of large language models for code?
    Through an online survey of 118 practitioners, we found that 81% of practitioners agreed that code readability is important. The key motivation is to reduce maintenance costs in the long term, while the key challenging factor is time constraints. Although code readability is currently improved via code review comments, 72% of practitioners agreed to consider adopting LLMs as an alternative. 39% of practitioners perceived that LLM-generated code is more readable than human-written code in general, followed by 34% perceived that the readability of both LLM-generated and human-written code is similar. These findings suggest that code readability is still critically important in the age of large language models.
  2. RQ2)
    How readable is the LLM-generated code compared to human-written code in practice?
    Through an examination of the readability of code generated by our LLM-powered code generation (HULA (takerngsaksiri2024human), powered by GPT-4 (openai2023gpt4)) and human-written code using our 144 internal Jira software development tasks, we found that the readability of LLM-generated code is comparable to human-written code (with a negligible to small effect size difference) based on various measurements (e.g., lines of code, code complexity, comment to code ratio). These findings demonstrate that the LLM-generated code can closely resemble human-written code in terms of readability.
Contributions. The paper contributions are as follows:
  • We conduct the first survey capturing practitioners’ perceptions of the current state of code readability and code readability of LLM-generated code, revealing generally positive feedback on code readability when adopting HULA in software development.
  • We revisit the importance and factors related to code readability from literature through the survey, aiming to raise awareness of code readability in practice.
  • We conduct the first empirical study of code readability comparison between human-written code and LLM-generated code in a real-world setting using our HULA framework, revealing promising results of the LLMs.
The results lead us to conclude that the readability of LLM-generated code is comparable to that of human-written code. While the survey reveals that the majority of practitioners perceive LLM-generated code as more readable or equally readable as human-written code, the empirical study supports the statements showing comparable results ranging from small to negligible effect sizes across multiple evaluation metrics.
To support the open science initiatives and increase the verifiability of our study, the survey and run results are available at https://github.com/awsm-research/CodeReadability-GenAI.
Paper Organization. Section LABEL:sec_background describes the background of HULA, our LLM-based software development agents framework. Section 2 presents the motivation and related work. Section 3 and Section 4 present the approach and results of RQ1 and RQ2, respectively. Section 5 discloses the threats to the validity of our study, while Section 6 draws the conclusion.
2. Motivation and Related Works
In this section, we discuss motivation and related work to formulate the research questions.
2.1. Motivation
Code readability is a cornerstone of software development at Atlassian, particularly as we integrate LLMs into Jira and other platforms (jiraai2024; takerngsaksiri2024human). With the adoption of LLM-powered tools, the size and complexity of codebases are expected to grow rapidly as the majority of the codebase could be generated by LLMs (adam2004). This paradigm shift raises concerns among Atlassian software engineers and customers regarding the readability of LLM-generated code, which may increase technical debts, incur maintenance costs, and negatively affect enterprise coding standards. Therefore, it is of utmost importance to understand whether readability retains its value in this LLM-assisted paradigm, and if so, why? And, how readable is LLM-generated code compared to human-written code?
2.2. Related Works
Code readability is a property that influences how easily a given piece of code can be read and understood (buse2008metric). Over the past 20 years (tashtoush2013impact; coleman2018aesthetics; fakhoury2018effect; mannan2018towards; alawad2019empirical; johnson2019empirical), prior studies have focused on investigating factors influencing code readability and developing metrics, models, and tools to quantify code readability. For example, Buse and Weimer (buse2008metric; buse2009learning) found that the average lines of code and average number of identifiers per line are closely associated with code readability. Similarly, Posnett et al. (posnett2011simpler) identified associations between readability and metrics such as the number of lines, Halstead Volume, and character entropy. Furthermore, Scalabrino et al. (scalabrino2016improving; scalabrino2018comprehensive) emphasized that readability is influenced not only by code structure but also by textual features such as code comments. Moreover, Alawad et al. (alawad2019empirical) observed a negative correlation between readability and code complexity, reporting that higher complexity is often associated with lower readability.
In the age of large language models (LLMs), studying code readability is more crucial than ever. While LLMs have now been seamlessly integrated into software development tools, workflows, and platforms like GitHub Copilot (copilot2024), IntelliJ IDEA (intellij2025), ChatGPT (openai2023gpt4), and Atlassian’s Jira AI (jiraai2024; takerngsaksiri2024human), they also introduce new challenges. Practitioners often raised concerns about the quality of LLM-generated code. For example, Liu et al. (liu2024refining) found that 47% of ChatGPT-generated code snippets suffer from maintainability issues. Majdinasab et al. (majdinasab2024assessing) discovered that 27% of code suggested by GitHub Copilot contains code vulnerabilities. Similarly, Perry et al. (perry2023users) found that users who had access to an AI assistant wrote significantly less secure code than those without access. Such concerns often lead to a lack of trust in the adoption of LLM-powered software development tools.
While many studies flag issues on LLM-generated code in the security aspect, it is still unclear on the readability aspect of how readable is LLM-generated code compared to human-written code. Recently, Madi (al2022readable) found that code generated by GitHub Copilot is comparable in complexity and readability to code written by human pair programmers in the live coding of a controlled environment. However, little is known about the code readability of LLM-generated code on production in real-world scenarios and how the practitioners perceived code readability in the age of large language models. The importance of code readability and the gap in literature lead us to formulate the following research questions:
  1. RQ1.
    How and why code readability is important in the age of large language models for code?
  2. RQ2.
    How readable is the LLM-generated code compared to human-written code in practice?
3. (RQ1) How and why code readability is important in the age of large language models for code?
3.1. Approach
The practitioners’ survey aims to explore four main topics of code readability: the importance, the challenges, the state of practice and the readability of LLM-generated code. In this section, we describe the design of our survey study and the participant selection process.
Table 1. Survey questions (excluding demographics).
Item Question
The Importance of Code Readability
Q1.1 How much of your time is devoted to reading code in your professional practice?
Q1.2 How important is code readability in your opinion?
Q1.3 What motivates you to prioritize code readability improvement in your professional practice?
The Challenges of Code Readability
Q2.1 How often do you prioritize code readability in your professional practice?
Q2.2 What factors prevent you from prioritizing the improvement of code readability in your professional practice?
The State of Code Readability
Q3.1 What are the factors you consider when improving the code readability?
Q3.2 Would you consider adopting Generative AI to improve code readability?
Q3.3 How do you improve your code readability in your professional practice?
Q3.4* If you improve code readability from automated tools, please kindly specify the name of the tools.
The Code Readability of Generative AI
Q4.1 How often do you use Generative AI for code generation/completion in your professional practice?
Q4.2 To what extent do you think model-generated code is readable compared to human code in professional practice?
$\dagger$ Likert scale of frequency Likert scale of agreement *Open-ended questions.
Survey. We design the 15-minute Google Form with four main sections: the importance of code readability, the challenges of code readability, the state of code readability, and the readability of LLM-generated code. The themes for the Likert scales of agreement questions (e.g., motivations for code readability) are derived from the literature review. Specifically, we categorized factors identified in the literature into themes, which formed the design of our survey questions (i.e., Q1.3, Q2.2, Q3.1). Excluding the demographic questions, the survey questions are shown in Table 1. The lists of derived themes are shown in Table 2 (Q1.3 and Q2.2) and Table 3 (Q3.1). Lastly, the survey obtained Ethical Permission from the Monash University Human Research Ethics Committee (MUHREC, Project ID 42299) before conducting the research.
Table 2. The Importance of Code Readability
Mentioned in the Literature.
No. Description Source 1 Bugs Preventing (piantadosi2022evolution; scalabrino2018comprehensive) 2 Code Complexity (fakhoury2019improving; tashtoush2023notional; alawad2019empirical; holst2021importance; fakhoury2018effect) 3 Code Comprehension (buse2009learning; posnett2011simpler; dorn2012general; scalabrino2018comprehensive; piantadosi2022evolution; sampaio2016software; fakhoury2018effect; johnson2019empirical) 4 Code Maintenance and Evolution (buse2008metric; buse2009learning; dorn2012general; scalabrino2016improving; scalabrino2018comprehensive; posnett2011simpler; sedano2016code; karanikiotis2020data; piantadosi2022evolution; sampaio2016software; lee2013study; piantadosi2020does; johnson2019empirical) 5 Code Standard & Aesthetic (lee2013study; borstler2016beauty; coleman2018aesthetics) 6 Design Quality (scalabrino2018comprehensive; fakhoury2019improving; mannan2018towards; tashtoush2013impact; piantadosi2022evolution; sampaio2016software) 7 Team Collaboration (holst2021importance; sampaio2016software; lee2013study; piantadosi2020does; johnson2019empirical)
Table 3. Coding Factors Related to Code Readability
Identified in the Literature.
No. Description Source 1 Code length (e.g., length of the identifier, line, function, file, etc.) (buse2008metric; buse2009learning; dorn2012general; tashtoush2013impact; posnett2011simpler) 2 Code structure/format (e.g., indentation, use of brackets, blank lines, etc.) (buse2008metric; buse2009learning; tashtoush2013impact; dorn2012general) 3 Code style/standard (e.g., PEP8, ESLint, Linter, etc.) (lee2013study; borstler2016beauty; coleman2018aesthetics; fakhoury2019improving) 4 Comment and documentation (buse2008metric; buse2009learning; dorn2012general; tashtoush2013impact; scalabrino2018comprehensive; scalabrino2016improving; karanikiotis2020data) 5 Descriptive naming variable and function (i.e., lexicon and semantics) (tashtoush2013impact; dorn2012general; scalabrino2018comprehensive; scalabrino2016improving; fakhoury2018effect; sedano2016code) 6 Function/file complexity (johnson2019empirical; fakhoury2018effect; tashtoush2023notional; alawad2019empirical; posnett2011simpler; sedano2016code; karanikiotis2020data)
Participant. We recruit participants through advertisements within Atlassian and public channels, including social media platforms like programming-related Facebook groups and personal LinkedIn networks. The participants can opt-in to win one of three gift cards of \$20 as a token of appreciation. Finally, during 3 weeks of advertisement in June 2024, we received 118 survey responses with diverse backgrounds. We summarize the attributes of the participants’ demographic in Figure 2.
3.2. Results
Table 4 to 7 show the results of this research question, covering four themes of code readability based on 118 practitioners’ responses.
The Importance: 81% of practitioners agree that code readability is important and the top motivation to improve code readability is to reduce maintenance cost and effort in the long term. Table 4 shows that 65% of practitioners answered spending more than half of their time reading code in their professional practice, while 81% of them were aware of the importance of code readability. Particularly, the top three motivations to prioritize code readability are to reduce maintenance costs and effort in the long term, to help the debugging and troubleshooting process, and to improve code comprehension for oneself and others.
Table 4. (RQ1) The Importance of Code Readability from Practitioners’ Perspective.
The Importance of Code Readability
Q1.1 How much of your time is devoted to reading code in your professional practice? 65%       11%
  Always (>75%)   Often (50-75%)   Sometimes (25-50%)   Rarely (<25%)   Never (0%)
Q1.2 How important is code readability in your opinion? 81%      7% Q1.3 What motivates you to prioritize code readability improvement in your professional practice? A. Improve code comprehension for oneself and others. 85%      8% B. Reduce maintenance cost and effort in the long term. 90%       5% C. Facilitate reusable and scalable development process. 76%       9% D. Enhance productivity and collaboration in the team. 81%      5% E. Reflect professionalism, aesthetic and industrial standard. 72%       8% F. Help debugging and troubleshooting process. 87%      6%   Strongly Agree   Agree   Neutral   Disagree   Strongly Disagree
Table 5. (RQ1) The Challenges of Code Readability from Practitioners’ Perspective.
The Challenges of Code Readability
Q2.1 How often do you prioritize code readability in your professional practice? 70%      4%
  Always (>75%)   Often (50-75%)   Sometimes (25-50%)   Rarely (< 25%)   Never (0%)
Q2.2 What factors prevent you from prioritizing the improvement of code readability in your professional practice? A. Time constraints due to project deadlines. 79%       9% B. Focus on code functionality over code readability. 65%       16% C. Focus on other performance metrics (e.g., speed and efficiency). 58%       17% D. Complexity of the codebase or legacy code makes the code readability improvement challenging. 70%       17% E. Resistance from team members or stakeholders. 61%       22% F. Limited resource or support on standards or best practices on code readability improvement. 63%       20%   Strongly Agree   Agree   Neutral   Disagree   Strongly Disagree
The Challenges: 79% of practitioners agree that the top challenging factor preventing them from prioritizing code readability is time constraints due to project deadlines. Table 5 shows that 70% of practitioners answered they always or often prioritize code readability in their professional practice. However, the top three factors that prevent them from prioritizing the improvement of code readability are the time constraints of project deadlines, the complexity of the codebase or legacy code, and the focus on code functionality over code readability.
Figure 2. The Demographic of Participants in RQ1 ($n$=118).
Table 6. (RQ1) The State of Code Readability from Practitioners’ Perspective.
The State of Code Readability
Q3.1 What are the factors you consider when improving the code readability?
A. Code length (e.g., length of identifier, line, function, file, etc.) 71%       11%
B. Code structure/format (e.g., indentation, use of brackets, etc.) 77%       7%
C. Code style/standard (e.g., PEP8, ESLint, Linter, etc.) 74%       11%
D. Comment and documentation 84%      4%
E. Descriptive naming variable/function 86%      4%
F. Function/file complexity 82%      5%
Q3.2 Would you consider adopting LLMs to improve code readability? 72%       8%
  Strongly Agree   Agree   Neutral   Disagree   Strongly Disagree
Q3.3 How do you improve your code readability in your professional practice? 68%       29%
  Manually by peer review comments   Manually self-improve code   Mixed   Static tools   AI-based tools Q3.4* If you improve code readability from automated tools, please kindly specify the name of the tools. A. ChatGPT   12 B. ESLint   5 C. Prettier   5 D. Copilot   3 E. KtLint   3 F. Black   2 *Open-ended questions.
The State of Practice: While manually improving code readability is still a dominant method with 68% of practitioners selection, 72% of practitioners agree to consider adopting LLMs to improve code readability. Table 6 shows that the most popular method for improving code readability in professional practice is still done manually by either peer review (41%) or self-improvement (27%). However, 17% of the practitioners have started to adopt LLMs to improve the code readability. The majority of them indicated that they used ChatGPT (n=12) as a tool. In fact, 72% of the practitioners are open to adopting LLMs to improve code. Additionally, the top three factors that practitioners consider when improving code readability are the descriptive naming variables and functions, the comments and documentation, and the complexity of functions or files.
Table 7. (RQ1) The Readability of LLM-generated Code from Practitioners’ Perspective.
The Readability of LLM-generated Code
Q4.1 How often do you use LLMs for code generation/completion in your professional practice? 54%       25%
  Always (>75%)   Often (50-75%)   Sometimes (25-50%)   Rarely (<25%)   Never (0%)
Q4.2 To what extent do you think LLM-generated code is readable compared to human code in professional practice? 27%     39%
  Human-written code is more readable   Both are similar   LLM-generated code is more readable
The Readability of LLM-generated Code: 39% of practitioners answer they perceive the LLM-generated code as more readable than the human-written code. Table 7 shows that 54% of practitioners answered using LLMs to generate or complete code more than half of their time in professional practice, indicating the current trend of high usage in LLM-generated code. Meanwhile, 39% of practitioners answered they perceive the LLM-generated code as more readable than the human-written code, followed by 34% of them answered both are similarly readable. The top justifications from the follow-up open-text question are logic complexity (n=25), code length (n=15), a number of loop conditions (n=12) and code structure/format (n=12). Specifically, a practitioner who voted LLM-generated code as more readable indicated that “Human code has too many logic conditions to check”. On the other hand, a practitioner who voted Human code as more readable indicates that “I find it easier to follow along.” Interestingly, some practitioners reason that “Actually, Human code is easier to understand in terms of solution, but having a lot of nested loop and conditions makes it harder to read and debug.”, indicates that understanding code and reading code can be as well on the different scales of decision.
4. (RQ2) How readable is the LLM-generated code compared to human-written code in practice?
4.1. Approach
The empirical study aims to understand code readability via static code analysis. In this section, we describe the design of our case study on HULA (takerngsaksiri2024human) over our internal dataset.
Figure 3. The Overview of the RQ2 Empirical Study Comparing Human-written Code to HULA-generated Code on Code Readability Measures.
Overview. Figure 3 shows the overview of the approach in this research question. Given a JIRA issue and a corresponding code repository, a software engineer writes a code, while AI coding agents, HULA, generate the codes to resolve the JIRA issue. Then, the human-written code and the HULA-generated code are assessed via eight code readability measurements. Lastly, we compare the results using the statistical analysis of the significance of Mann-Whitney U Test (mann1947test) and the effect size of Cliff’s Delta (cliff1993dominance).
Large Language Model. We evaluate the code readability using our Human-in-the-loop Software Development Agents framework, HULA (takerngsaksiri2024human). HULA is Atlassian’s LLM-based agents framework for software development available internally in Jira Software. Specifically, the framework is used to generate a code patch given a code repository and a description of the Jira issue to be resolved. In this version, HULA uses GPT-4 (openai2023gpt4) as a based LLM and the framework is evaluated without additional human feedback.
Datasets. We evaluate HULA via the internal dataset (takerngsaksiri2024human), which includes a set of completed software development tasks (i.e., Jira issues, repositories and corresponding pull requests from BitBucket) at Atlassian. The dataset consists of 144 Jira issues with 250 coding files spanning six programming languages: TypeScript, Java, Kotlin, Python, Go and Scala. The distribution of programming languages is illustrated in Figure 4.
Figure 4. The Distribution of Programming Languages in the Dataset (n=250).
Code Readability Measurement. Following the recent study on code readability (al2022readable), we cautiously select the following eight static analysis metrics that are language-agnostic (i.e., can be used in multilingual) via the Multimetric library(multimetric2024).
First, we use Line of Code to evaluate the total number of lines of code in the file, expressing the file size. The Comment Ratio is used to evaluate the percentage of the lines of comment to code, expressing the descriptive of the code. Then, Cyclomatic Complexity (mccabe1976complexity) is used to evaluate the number of decision paths of the code, expressing the code complexity. Next, Maintainability Index (Oman1992maintainability) is used to measure how maintainable the code is. Lastly, we utilize Halstead Metrics (halstead1977elements), a set of metrics designed to quantify various aspects of code. These metrics are employed to evaluate key attributes, including Difficulty (the difficulty of the program to write or understand), Vocabulary (the number of unique operators and unique operands occurrence), Volume (the program length and vocabulary size), and Time Required (the estimated time for implementation).
Below, we list the detailed calculations of each metric.
Figure 5. (RQ2) The Scores of Code Readability Measures between HULA-generated and Human-written Code.
Figure 6. (RQ2) The Median Scores of Code Readability Measures between Human-written and HULA-generated code by Programming Languages.
  1. (1)
    Line of Code is the total number of lines of code.
  2. (2)
    Comment Ratio is the percentage of the lines of comment to the lines of code.
  3. (3)
    Cyclomatic Complexity (mccabe1976complexity) or McCabe number is the number of decisions path of code according. Specifically, the complexity is defined as $C=E-N+2$, where $E$ is the number of edges on the control flow graph and $N$ is the number of nodes on the control flow graph.
  4. (4)
    Maintainability Index (Oman1992maintainability) is a metric to measure how maintainable the code is. We use the Microsoft formula (microsoft2022maintainability) defined as $Max(0,(171-5.2*ln(HalsteadVolume)-0.23*(McCabe)-16.2*ln(LoC))*100/171)$.
  5. (5)
    Halstead Difficulty (halstead1977elements) defines as $D=(n1/2)*(N2/n2)$, where $n1$ is the number of distinct operators, $n2$ is the number of distinct operands, $N1$ is the total number of operators, and $N2$ is the total number of operands.
  6. (6)
    Halstead Vocabulary (halstead1977elements) defines as $n=n1+n2$.
  7. (7)
    Halstead Volume (halstead1977elements) defines as $V=(N1+N2)*log_{2}(n)$.
  8. (8)
    Halstead Time Required (halstead1977elements) defines as $T=(D*V)/18$
4.2. Results
Figure 5 shows the results of this research question, covering the case study of HULA on our internal dataset. Table 8 shows the statistical significance and the effect size of the result. Lastly, Figure 6 shows the detailed analysis of code readability measurement by programming languages.
The generated code is comparable to the human-written code with a negligible to small statistical difference. Figure 5 shows that HULA, with a GPT-4 based model, generates code that is highly similar to humans with only a slightly longer code but with no statistically significant difference in Cyclomatic Complexity and comment ratio. On Halstead’s metrics, our LLM-generated code has marginally higher values on difficulty, vocabulary, volume, and time required. These results indicate that the HULA framework can produce code that closely resembles human-written code in terms of code readability, indicating the promising performance of LLM in production.
The HULA-generated code is comparable to the human-written code specifically in Java, Kotlin, Go and Scala programming languages. Figure 6 shows the static analysis results by programming languages of code generated by Hula and written by humans. The case study on the Hula framework reveals two different patterns across the six studied programming languages. First, in Java, Kotlin, Go and Scala programming languages, HULA-generated code shows comparable results across all code readability measures to that written by humans. Second, in TypeScript and Python programming languages, the Hula-generated code tends to have longer lines of code and slightly lower maintenance scores than the human-written code by small effect sizes.
Table 8. (RQ2) The mean significance from Mann-Whitney U Test (mann1947test) and Effect Size measured by the Cliff’s Delta (cliff1993dominance) based on the comparison of HULA and Human code.
Metric P-Value Effect Size
Line of Code <0.001 0.230 (small)
Comment Ratio 0.005 0.124 (negligible)
Cyclomatic Complexity 0.622 0.015 (negligible)
Maintainability Index <0.001 -0.225 (small)
Halstead Difficulty <0.001 0.201 (small)
Halstead Vocabulary <0.001 0.202 (small)
Halstead Volume <0.001 0.208 (small)
Halstead Time <0.001 0.205 (small)
5. Threat to Validity
In this section, we disclose the threats to the validity.
Threats to internal validity. We use an online survey to study practitioners’ perceptions of code readability. We design the questions based on literature reviews to mitigate the hallucination on factors related to code readability. However, we acknowledge the limitations that the justifications might be beyond the literature review which we sourced from.
Threats to external validity. Our survey was derived from the perceptions of 118 practitioners with different backgrounds and work experiences in coding (see Figure 2). Nonetheless, we acknowledge that the results may differ when surveying different population groups. Additionally, the empirical study is a case study of our HULA framework, we acknowledge that the results may differ when experimenting with different models and datasets.
Additionally, the empirical study was conducted on code generation tasks using GPT-4 as a based model and the Atlassian internal dataset. Efforts were made to ensure diversity in programming languages and task contexts. However, it is acknowledged that the findings may vary when applied to different datasets. Therefore, future research is encouraged to investigate code readability across a broader range of contexts.
Lastly, while the HULA framework typically operates with a human-in-the-loop—where software engineers provide feedback to guide the outputs of the LLM-based agent, in this study, all human feedback and intervention were intentionally excluded. The effort was made to avoid introducing additional information to the LLM-based agents and to ensure a fair comparison between code written by humans and code generated by LLM-based agents. Future studies could compare results from both autonomous and human-in-the-loop settings to better understand their respective trade-offs.
Threat to construct validity. We use an empirical study with static code analysis as a proxy for code readability measurement between LLM-generated and Human code. We cautiously follow the existing work on selecting the metrics representing code readability. However, we acknowledge the limitations that the metrics might not completely represent the code readability from all perspectives.
Last but not least, although we may not be able to disclose the source code of the evaluation dataset due to industrial confidentiality reasons, we are confident that the evaluation offers valuable insight into the practical use of LLMs in real-world development workflows and reflects industrial standards for production-level code, paving the way forward the future of AI-powered software development tools.
6. Conclusion and Implications
In this paper, we investigate the practitioners’ perceptions on the importance, the challenges, and the state of practice of code readability, and investigate the readability of LLM-generated code and human-written code in the context of enterprise software development tasks. Our findings underscore that readability remains a critical aspect of software development, even in the age of large language models (LLMs). Moreover, the code generated by our LLM-powered framework, HULA, is shown to be comparable in readability to human-written code. This supports the establishment of appropriate trust and drives the broad adoption of our LLM-powered software development platform. Practitioners can confidently integrate LLMs into their workflows, knowing that code readability will not be compromised. This fosters more effective teamwork, simplifies maintenance tasks, and ensures the long-term success of software projects.